Every morning, Don Miguel exercises by walking a certain distance at an average of 4 kph, rests for 20 min. and rides back along the same road on his chauffer-driven limousine at an average speed of 60 kph. If the total time of travel is one hr. and 15 min. how far does he walk
Answer by stanbon(75887)

Every morning, Don Miguel exercises by walking a certain distance at an average of 4 kph, rests for 20 min. and rides back along the same road on his chauffer-driven limousine at an average speed of 60 kph. If the total time of travel is one hr. and 15 min. how far does he walk
--------------
Walking DATA:
rate = 4kph; distance = x km ; time = x/4 hrs
------
Riding DATA:
rate = 60 kph; distance = x km ; time = x/60 hrs
----------
Equation:
time + time + time = total time
x/4 + x/60 + 1/3 = 5/4 hr
----
Multiply thru by 60 to get:
15x + x + 20 = 75
16x = 55
x = 55/16 = 3.4375 km (distance he walked)
========================
Cheers,
Stan H.